# Learning General Representation of 12-Lead Electrocardiogram With a Joint-Embedding Predictive Architecture

### v2

@@ Abstract
We propose a self-supervised learning method for 12-lead Electrocardiogram (ECG) analysis, named ECG Joint Embedding Predictive Architecture (ECG-JEPA). ECG-JEPA employs a masking strategy to learn semantic representations of ECG data. Unlike existing methods, ECG-JEPA predicts at the hidden representation level rather than reconstructing raw data. This approach offers several advantages in the ECG domain: (1) it avoids producing unnecessary details, such as noise, which is common in standard ECG; and (2) it addresses the limitations of naïve L2 loss between raw signals. Another key contribution is the introduction of a special masked attention tailored for 12-lead ECG data, Cross-Pattern Attention (CroPA). CroPA enables the model to effectively capture inter-patch relationships. Additionally, ECG-JEPA is highly scalable, allowing efficient training on large datasets. Our code is openly available at https://github.com/sehunfromdaegu/ECG_JEPA.
@@ 1 Introduction
Electrocardiograms (ECG) provide a non-invasive method to measure the electrical activity of the heart over time, serving as a crucial tool for diagnosing various cardiac conditions. While numerous supervised methods have been developed to detect heart diseases using ECG data [1, 2, 3], these models often face significant performance degradation when applied to data distributions different from those on which they were trained.
Self-supervised learning (SSL) has emerged as a powerful paradigm for learning general representations across various domains, including natural language processing (NLP) [4, 5, 6], computer vision (CV) [7, 8, 9], and video analysis [10, 11]. Despite its promise, applying SSL to ECG data presents unique challenges. For instance, data augmentation, which is essential in many SSL architectures, is more complex for ECG than for computer vision data. Simple transformations like rotation, scaling, and flipping, effective in CV, can distort the physiological meaning of ECG signals. Additionally, ECG recordings often contain artifacts and noise, which cause autoencoder-based SSL models to struggle with reconstructing raw signals. These architectures may also miss visually subtle but diagnostically critical features, such as P-waves and T-waves, which are imperative for diagnosing certain cardiac conditions.
In this work, we propose a novel ECG Joint-Embedding Predictive Architecture (ECG-JEPA) tailored for 12-lead ECG data, effectively addressing the aforementioned challenges. ECG-JEPA utilizes a transformer architecture to capture the semantic meanings of ECG. By masking several patches of the ECG data, ECG-JEPA predicts the abstract representations of the missing patches, indicating a high-level understanding of the data. Additionally, we develop a novel masked-attention for multi-lead ECG data, coined Cross-Pattern Attention (CroPA). CroPA incorporates clinical knowledge into the model as an inductive bias, guiding it to focus on clinically relevant patterns and relationships across leads.
We demonstrate the following contributions of ECG-JEPA:
@@
  • ECG-JEPA can also recover important ECG features, including heart rate and QRS duration, which are classical indicators used in ECG evaluation. This is the first work to demonstrate that learned representations can effectively recover various ECG features.
  • ECG-JEPA is highly scalable, allowing efficient training on large datasets. For instance, ECG-JEPA is trained for only 100 epochs, yet outperforms other ECG SSL models on most downstream tasks, taking approximately 26 hours on a single RTX 3090 GPU.
Figure 1: 12-lead ECG. The standard 12-lead ECG captures electrical activity from different angles around the heart, providing comprehensive cardiac assessment. Artifacts and noise, such as powerline interference and baseline wander, are common in ECG recordings. The figure shows an example of a 12-lead ECG with baseline wander.
@@ 2 Background
Self-Supervised Learning (SSL) facilitates learning abstract representations from input data without the need for labeled data, which is particularly beneficial in medical domains where labeled data is scarce and expensive. SSL leverages inherent data patterns to learn useful representations, allowing models to adapt to various downstream tasks with greater robustness to data imbalances [12]. In this section, we briefly explain key SSL techniques in Sections 2.1 and 2.2, including generative architectures and joint-embedding predictive architectures. Additionally, we provide an overview of the ECG and its key features in Sections 2.3 and 2.4, highlighting the critical characteristics essential for understanding ECG data.
2.1 Generative Architectures
Generative architectures involve reconstructing an input $x$ from its degraded version $x^{\prime}$ using an encoder-decoder framework. The idea behind generative architectures is that the ability to reconstruct the clean data from a corrupted one indicates a good understanding of the data. The encoder maps the perturbed input $x^{\prime}$ into a latent representation, which the decoder then uses to generate the original input $x$ [13].
2.2 Joint-Embedding Predictive Architectures
Joint-Embedding Predictive Architectures (JEPA) [14] process pairs $x$ and their corrupted versions $x^{\prime}$ to obtain representations $z$ and $z^{\prime}$ through encoders. Unlike generative architectures that predict in the input space, JEPA predicts in the hidden representation space by reconstructing $z$ from $z^{\prime}$. This approach effectively avoids the challenge of predicting unpredictable details, which is common in biological signals.
@@ 2.3 Electrocardiogram (ECG)
The electrocardiogram (ECG) is a non-invasive diagnostic method that records the heart’s electrical activity over time using electrodes placed on the skin. The standard 12-lead ECG captures electrical activity from multiple angles. These 12 leads are categorized into limb leads (I, II, III), augmented limb leads (aVR, aVL, aVF), and chest leads (V1-V6). Each lead provides unique information about the heart’s electrical activity, offering a comprehensive view that aids in diagnosing various cardiac conditions. Refer to Figure 1 for an illustration.
In this work, we use only 8 leads (I, II, V1-V6) as the remaining 4 leads (III, aVR, aVL, aVF) can be derived from linear combinations of the 8 leads (see C). This choice maintains the necessary diagnostic information while optimizing computational efficiency.
@@
2.4 ECG Features
[…]
Figure 3: ECG-JEPA. The teacher network processes the complete ECG data, generating fully contextualized representations. In contrast, the student network receives only the non-masked patches. The student’s representations are then passed to the predictor, which reconstructs the fully contextualized representations produced by the teacher. The predicted 6 representations are then compared with the corresponding teacher’s representation. This figure illustrates $L=3$ leads with $N=5$ subintervals and $Q=3$ unmasked subintervals.
@@ 3.3 Cross-Pattern Attention (CroPA)
4 Related Work
This section provides an overview of existing SSL methods that have been applied to ECG data. We compare ECG-JEPA with these methods on various downstream tasks in Section 6.
4.1 Contrastive Learning
Contrastive learning has been widely explored in various research works for its effectiveness in unsupervised representation learning. Contrastive Predictive Coding (CPC) [15] aims to capture useful data representations by predicting future observations in the latent space. It uses sequential data where the “true future" is treated as the positive sample, and other observations within the same sequence are treated as negative samples. SimCLR [7] and MoCo [16] also leverage contrastive approaches, with positive samples being augmented versions of the same image and negative samples being other images. SimCLR relies on large batch sizes and extensive data augmentation, while MoCo uses a momentum encoder to maintain a dynamic dictionary of encoded samples, enabling more stable training with smaller batch sizes.
4.2 Generative Architectures
Generative architectures, first popularized by BERT [4] in natural language processing (NLP), have since inspired a variety of approaches in other domains. For example, in computer vision, Masked Autoencoders (MAE) [8] extend this concept by masking portions of an input image and training the model to reconstruct the missing parts, similar to how BERT predicts masked tokens in text. Recently, the authors of [17] pointed out that generative architectures often prioritize learning the principal subspaces of the data, which may limit their effectiveness in capturing semantic representations for perceptual tasks.
Figure 4: Cross-Pattern Attention (CroPA). The patch in the middle attends only to the colored patches, which are in the same lead and temporal space.
4.3 Joint-Embedding Predictive Architectures
In contrast to generative approaches, which focus on reconstructing data in the input space, joint-embedding predictive architectures (JEPA) aim to predict in the representation space, allowing models to learn more abstract and semantic features. An example of this approach is Data2vec [18], which introduced an SSL architecture that predicts the average of the encoder’s final layer outputs. Other examples include I-JEPA [9] and V-JEPA [11], which predict the hidden representations of masked patches, enabling semantic representation learning for images and videos, respectively.
4.4 Self-Supervised Learning for ECG
Various SSL models have been explored for 12-lead ECG data. Contrastive Multi-segment Coding (CMSC) [19] splits an ECG into two segments, encouraging similar representations for compatible segments while separating incompatible ones. Contrastive Predictive Coding (CPC) [15], applied in [20], predicts future ECG representations, but its reliance on LSTM modules makes it inefficient for large datasets. More recently, [21] introduced masked autoencoders for ECG, proposing temporal and channel masking strategies, Masked Time Autoencoder (MTAE) and Masked Lead Autoencoder (MLAE). Similarly, [22] proposed ST-MEM, which masks random time intervals for each lead. However, both MLAE and ST-MEM may struggle with the high correlations between ECG leads, making some predictions too easy.
@@ 6 Experiments
In most cases, ECG-JEPA consistently outperforms other SSL methods that rely on hand-crafted augmentations, highlighting its effectiveness in learning generalizable representations from ECG data. In our experiments, E-JEPArb and E-JEPAmb refer to ECG-JEPA models trained using random masking and multi-block masking strategies, respectively.
Table 1: Linear evaluation on multi-label task.
PTB-XL CPSC2018
Method Epochs AUC F1 AUC F1
ST-MEM 800 0.896 0.662 0.964 0.752
SimCLR 300 0.866 0.624 0.890 0.523
CMSC 300 0.802 0.472 0.767 0.206
CPC 100 0.620 0.167 0.687 0.091
E-JEPArb 100 0.906 0.690 0.969 0.769
E-JEPAmb 100 0.912 0.712 0.971 0.789
Table 2: Linear evaluation on multi-class task.
PTB-XL CPSC2018
Method Epochs AUC F1 AUC F1
MoCo v3111Score as reported in [22]. 800 0.739 0.142 0.712 0.080
MTAE111 800 0.807 0.437 0.818 0.349
MLAE111 800 0.779 0.382 0.794 0.263
ST-MEM 800 0.888 0.566 0.973 0.805
SimCLR 300 0.842 0.496 0.918 0.624
CMSC 300 0.796 0.442 0.787 0.391
CPC 100 0.600 0.201 0.672 0.210
E-JEPArb 100 0.894 0.616 0.974 0.805
E-JEPAmb 100 0.896 0.628 0.973 0.819
@@ 6.1 Linear Evaluation
Tables 1 and 2 present the results of our linear evaluation on the PTB-XL and CPSC2018 datasets. We train a linear classifier on top of the frozen representations for 10 epochs and evaluate its performance on downstream tasks. Further training beyond 10 epochs does not lead to any significant improvement in performance. As shown in the tables, ECG-JEPA consistently outperforms other SSL methods, demonstrating superior efficiency and effectiveness with substantially reduced computational resources.
@@ 6.2 Fine-tuning
Fine-tuning is another method to evaluate the quality of learned representations, as it tests the model’s ability to adapt its pre-trained features to new tasks. In the fine-tuning process, we add a linear classification head at the end of the encoder and train the entire network for 10 epochs. Similar to linear evaluation, training for 10 epochs is sufficient, as further training does not lead to additional performance gains. Fine-tuning can potentially enhance performance beyond what is achieved with linear evaluation alone.
Table 3 presents the results of fine-tuning on the PTB-XL and CPSC2018 datasets. ECG-JEPA is compared with other SSL methods as well as supervised methods in a multi-class classification setting, where the student network is trained directly for the supervised task. The results indicate that ECG-JEPA achieves the highest AUC and F1 scores on PTB-XL and the highest AUC on CPSC2018. These findings underscore ECG-JEPA’s effectiveness in fine-tuning scenarios, demonstrating its ability to refine and adapt learned representations for improved performance in specific classification tasks.
@@
PTB-XL CPSC2018
Method Epochs AUC F1 AUC F1
@@
MoCo v3111 800 0.913 0.644 0.967 0.838
MTAE111 800 0.910 0.613 0.961 0.769
MLAE111 800 0.915 0.625 0.973 0.816
CMSC111 800 0.877 0.510 0.938 0.717
@@
CPC [^2] We did not fine-tune CPC due to its slow training process. 100 - - - -
E-JEPArb 100 0.944 0.710 0.980 0.821
E-JEPAmb 100 0.937 0.680 0.983 0.799
6.3 Low-shot Linear Evaluation
Table 4: Low-shot linear evaluation on the multi-label PTB-XL task. The mean and standard deviation of macro AUC are reported for 1% (192 samples) and 10% (1923 samples) of the training set, selected three times independently.
@@ 6.4 Reduced Lead Evaluation
Table 5: Reduced lead evaluation. Linear evaluation of PTB-XL multi-label classification in single-leade (II) and dual-lead (II and V1) configurations.
@@ 6.5 ECG Feature Extraction
Extracting ECG features is crucial for diagnosing and monitoring cardiac conditions. In this experiment, we assess the model’s ability to extract key features such as heart rate and QRS duration from the learned representations of the PTB-XL dataset. Unlike classification tasks, which focus on perceptual patterns, ECG features are directly tied to the signal’s morphology. For example, heart rate is derived from the average distance between R-peaks, while QRS duration is measured by the average width of the QRS complex.
@@
Table 6: ECG feature extraction. Means and standard deviations of absolute differences between the predicted and extracted values for heart rate and QRS duration. Small values indicate better performance. The final row provides the overall mean and standard deviation for both features across all samples in the PTB-XL test set.

### v4

@@ Abstract
Electrocardiogram (ECG) captures the heart’s electrical signals, offering valuable information for diagnosing cardiac conditions. However, the scarcity of labeled data makes it challenging to fully leverage supervised learning in medical domain. Self-supervised learning (SSL) offers a promising solution, enabling models to learn from unlabeled data and uncover meaningful patterns. In this paper, we show that masked modeling in the latent space can be a powerful alternative to existing self-supervised methods in the ECG domain. We introduce ECG-JEPA, a SSL model for 12-lead ECG analysis that learns semantic representations of ECG data by predicting in the hidden latent space, bypassing the need to reconstruct raw signals. This approach offers several advantages in the ECG domain: (1) it avoids producing unnecessary details, such as noise, which is common in ECG; and (2) it addresses the limitations of naïve L2 loss between raw signals. Another key contribution is the introduction of Cross-Pattern Attention (CroPA), a specialized masked attention mechanism tailored for 12-lead ECG data. ECG-JEPA is trained on the union of several open ECG datasets, totaling approximately 180,000 samples, and achieves state-of-the-art performance in various downstream tasks including ECG classification and feature prediction. Our code is openly available at https://github.com/sehunfromdaegu/ECG_JEPA.
@@ 1 Introduction
Electrocardiography is a non-invasive method to measure the electrical activity of the heart over time, serving as a crucial tool for diagnosing various cardiac conditions. While numerous supervised methods have been developed to detect heart diseases using ECG data [1, 2, 3], these models often face significant performance degradation when applied to data distributions different from those on which they were trained. This challenge points to the need for more flexible approaches that can learn robust, transferable representations from ECG data.
Self-supervised learning (SSL) offers an alternative approach by learning general representations in diverse domains, such as natural language processing (NLP) [4, 5, 6], computer vision (CV) [7, 8, 9], and video analysis [10, 11]. Despite this promise, the application of SSL to ECG data presents unique challenges. For instance, data augmentation, which is essential in many SSL architectures, is more complex for ECG than for computer vision data. Simple transformations like rotation, scaling, and flipping, effective in CV, can distort the physiological meaning of ECG signals. Additionally, ECG recordings often contain artifacts and noise, which cause autoencoder-based SSL models to struggle with reconstructing raw signals. These architectures may also miss visually subtle but diagnostically critical features, such as P-waves and T-waves, which are imperative for diagnosing certain cardiac conditions.
In this work, we propose ECG Joint-Embedding Predictive Architecture (ECG-JEPA) tailored for 12-lead ECG data, effectively addressing the aforementioned challenges. ECG-JEPA utilizes a transformer architecture to capture the semantic meaning of the ECG. By masking several patches of the ECG, ECG-JEPA predicts abstract representations of the missing segments, indicating a high-level understanding of the data. Additionally, we develop a novel masked-attention for multi-lead ECG data, chich we call Cross-Pattern Attention (CroPA). CroPA incorporates clinical knowledge into the model as an inductive bias, guiding it to focus on clinically relevant patterns and relationships across leads.
Figure 1: 12-lead ECG with baseline wander artifact.
Our contributions are as follows:
@@
  • ECG-JEPA can also recover important ECG features, including heart rate and QRS duration, which are classical indicators used in ECG evaluation. This is the first work to demonstrate that learned representations can effectively recover ECG features.
  • ECG-JEPA is highly scalable, allowing efficient training on large datasets. For instance, ECG-JEPA is trained for only 100 epochs, yet outperforms other ECG SSL models on most downstream tasks, taking approximately 22 hours on a single RTX 3090 GPU.
In summary, ECG-JEPA introduces a robust SSL framework for 12-lead ECG analysis, overcoming traditional SSL limitations with clinically inspired design elements, scalable architecture, and demonstrated effectiveness on a wide range of tasks.
@@ 2 Background
Self-Supervised Learning (SSL) facilitates learning abstract representations from input data without the need for labeled data, which is particularly beneficial in medical domains where labeled data is scarce and expensive. SSL leverages inherent data patterns to learn useful representations, allowing models to adapt to various downstream tasks with greater robustness to data imbalances [12]. We begin in Section 2.1 with an overview of the ECG and its key features, highlighting the critical characteristics essential for understanding ECG data. In Sections 2.2 and 2.3, we briefly explain key SSL techniques and their specific applications to ECG, respectively.
@@ 2.1 Electrocardiogram (ECG)
The electrocardiogram (ECG) is a non-invasive diagnostic method that records the heart’s electrical activity over time using electrodes placed on the skin. The standard 12-lead ECG captures electrical activity of the heart from multiple angles. These 12 leads are categorized into limb leads (I, II, III), augmented limb leads (aVR, aVL, aVF), and chest leads (V1-V6). Each lead provides unique information about the heart’s electrical activity, offering a comprehensive view that aids in diagnosing various cardiac conditions. Refer to Figure 1 for an illustration.
[…]
In this work, we use only 8 leads (I, II, V1-V6) as the remaining 4 leads (III, aVR, aVL, aVF) can be derived from linear combinations of the 8 leads following the Einthoven’s law [13]:
[…]
$\text{aVR}=-(\text{I}+\text{II})/2$
$\text{aVL}=(\text{I}-\text{II})/2$
$\text{aVF}=(\text{II}-\text{I})/2.$
This choice maintains the necessary diagnostic information while optimizing computational efficiency.
@@
Figure 3: ECG-JEPA training overview. For illustration, we use $L=3$, $N=5$ subintervals and $Q=3$ unmasked subintervals.
2.2 Self-Supervised Learning Architectures
Self-supervised learning can be broadly categorized into contrastive and non-contrastive methods. Non-contrastive methods can be further divided into generative and non-generative architectures. For a broader introduction to SSL, see [14].
In contrastive learning, the model is encouraged to produce similar representations for semantically related inputs $x^{{}^{\prime}}$ and $x^{{}^{\prime\prime}}$, while pushing apart the representations of unrelated inputs $x^{{}^{\prime}}$ and $y^{{}^{\prime}}$. SimCLR [7] is one of the most popular contrastive methods, using two different augmentations of a single input $x$ to form semantically similar pairs $x^{{}^{\prime}}$ and $x^{{}^{\prime\prime}}$.
Beyond contrastive methods, generative architectures have been particularly successful in recent large language models [4, 5, 6] and in computer vision [8]. Generative architectures involve reconstructing a sample $x$ from its degraded version $x^{\prime}$ using an encoder-decoder framework. The premise is that reconstructing clean data from a corrupted version reflects the model’s deep understanding of the underlying data structure. The encoder maps the perturbed input $x^{\prime}$ into a latent representation, which the decoder then uses to reconstruct the original input $x$ [15]. Recently, the authors of [16] observed that generative architectures prioritize learning principal subspaces of the data, which may limit their capacity to capture semantic representations for perceptual tasks.
As an alternative, non-generative methods have shown promise across domains, including computer vision [17, 18, 19, 9] and video analysis [11]. Among these, the Joint-Embedding Predictive Architecture (JEPA) [20] processes an input pair $x$ and its corrupted versions $x^{\prime}$ to obtain representations $z$ and $z^{\prime}$ through encoders. Unlike generative architectures that make predictions in the input space, JEPA performs prediction in the latent space by reconstructing $z$ from $z^{\prime}$. This approach effectively avoids the challenge of predicting unpredictable details, a common issue in biological signals.
2.3 Related Works
Several studies have worked on capturing semantically meaningful representations of 12-lead ECG data. Contrastive Multi-segment Coding (CMSC) [21] splits an ECG into two segments, encouraging similar representations for compatible segments while separating incompatible ones. Contrastive Predictive Coding (CPC) [22], applied in [23], predicts future ECG representations in a contrastive manner, but its reliance on LSTM modules makes it inefficient for large datasets. More recently, [24] introduced masked autoencoders for ECG, proposing temporal and channel masking strategies, Masked Time Autoencoder (MTAE) and Masked Lead Autoencoder (MLAE). Similarly, [25] proposed ST-MEM, which masks random time intervals for each lead. However, both MLAE and ST-MEM may struggle with the high correlations between ECG leads, potentially oversimplifying the prediction task.
@@ 3.3 Cross-Pattern Attention (CroPA)
Figure 4: Cross-Pattern Attention (CroPA). The patch in the middle attends only to the colored patches.
Figure 5: Squares following the encoder represent the representations of ECG patches. The representations are subsequently averaged through a pooling layer, with the resulting vector (highlighted in cyan) serving as an abstract representation of the ECG data.
3.4 ECG representation
After training, we use only the student network as the encoder. The encoder outputs are average-pooled to obtain the final ECG representation, which serves as the feature vector for downstream tasks. See Figure 5 for an illustration.
@@ 5 Experiments
Table 1: Linear evaluation on multi-label and multi-class tasks. Our proposed method outperforms all baselines, achieving the highest AUC and F1 scores across both tasks and datasets.
Multi-label Task Multi-class Task
Method Epochs PTB-XL CPSC2018 PTB-XL CPSC2018
AUC F1 AUC F1 AUC F1 AUC F1
ST-MEM 800 0.896 0.662 0.964 0.752 0.888 0.566 0.973 0.805
SimCLR 300 0.866 0.624 0.890 0.523 0.842 0.496 0.918 0.624
CMSC 300 0.802 0.472 0.767 0.206 0.796 0.442 0.787 0.391
CPC 100 0.620 0.167 0.687 0.091 0.600 0.201 0.672 0.210
MoCo v31 800 - - - - 0.739 0.142 0.712 0.080
MTAE1 800 - - - - 0.807 0.437 0.818 0.349
MLAE1 800 - - - - 0.779 0.382 0.794 0.263
ECG-JEPArb 100 0.906 0.690 0.969 0.769 0.894 0.616 0.974 0.805
ECG-JEPAmb 100 0.912 0.712 0.971 0.789 0.896 0.628 0.973 0.819
  • 1
    Scores reported in [25]; results for multi-label tasks were not available.
@@
In most cases, ECG-JEPA consistently outperforms other SSL methods that rely on hand-crafted augmentations, highlighting its effectiveness in learning generalizable representations. In our experiments, ECG-JEPArb and ECG-JEPAmb refer to ECG-JEPA models trained using random masking and multi-block masking strategies, respectively.
@@ 5.1 Linear Evaluation
Table 1 present the results of our linear evaluation on the PTB-XL and CPSC2018 datasets. We train a linear classifier on top of the frozen representations for 10 epochs and evaluate its performance on downstream tasks. Further training beyond 10 epochs does not lead to any significant improvement in performance. As shown in the tables, ECG-JEPA consistently outperforms other SSL methods, demonstrating superior efficiency and effectiveness with substantially reduced computational resources.
@@ 5.2 Fine-tuning
Fine-tuning is another method to evaluate the quality of learned representations, as it tests the model’s ability to adapt its pre-trained features to new tasks. We add a linear classification head at the end of the encoder and train the entire network for 10 epochs. Similar to linear evaluation, training for 10 epochs is sufficient, as further training does not lead to additional performance gains. Fine-tuning can potentially enhance performance beyond what is achieved with linear evaluation alone.
Table 2 presents the results of fine-tuning on the PTB-XL and CPSC2018 datasets. ECG-JEPA is compared with other SSL methods as well as supervised methods in a multi-class classification setting, where the student network is trained directly from the scratch. The results indicate that ECG-JEPA achieves the highest AUC and F1 scores on PTB-XL and the highest AUC on CPSC2018.
5.3 Low-shot Linear Evaluation
@@
Method Epochs PTB-XL CPSC2018
AUC F1 AUC F1
@@
MoCo v31 800 0.913 0.644 0.967 0.838
MTAE1 800 0.910 0.613 0.961 0.769
MLAE1 800 0.915 0.625 0.973 0.816
CMSC1 800 0.877 0.510 0.938 0.717
@@
CPC2 100 - - - -
ECG-JEPArb 100 0.944 0.710 0.980 0.821
ECG-JEPAmb 100 0.937 0.680 0.983 0.799
  • 1
    Scores reported in [25].
  • 2
    We did not fine-tune CPC due to its slow training process.
Table 3: Low-shot linear evaluation on the multi-label PTB-XL. The mean and standard deviation of macro AUCs are reported for 1% (192 samples) and 10% (1923 samples) of the training set, selected three times independently.
@@ 5.4 Reduced Lead Evaluation
Table 4: Reduced lead evaluation. Linear evaluation of PTB-XL multi-label classification in single-leade (II) and dual-lead (II and V1).
@@ 5.5 ECG Feature Extraction
Extracting ECG features is crucial for diagnosing and monitoring cardiac conditions. In this experiment, we assess the model’s ability to extract key features such as heart rate and QRS duration from the learned representations of the PTB-XL dataset. Unlike classification tasks, which focus on perceptual patterns, ECG features are directly tied to the signal’s morphology.
@@
Table 5: ECG feature prediction results on PTB-XL multi-lable test set. The mean heart rate and QRS duration in the test set are 70.01 BPM (±$\pm 17.65$) and 90.48 ms (±$\pm 17.02$), respectively.
Mean Absolute Error